OpenStudy (anonymous):

A certain part if the electromagnetic spectrum ranges from 350nm to 820nm . What is the highest frequency associated with this portion of spectrum?

OpenStudy (anonymous):

use E = h f = h c /lambda with wavelength lambda in m, E in J, c as d x 10^8 m/s

OpenStudy (anonymous):

f is frequency

OpenStudy (anonymous):

what are E and h?

OpenStudy (anonymous):

Electric field?

OpenStudy (anonymous):

^ that doesn't make sense, I know.

OpenStudy (anonymous):

E =energy of photon h= Planck's constant

OpenStudy (anonymous):

In the plank's constant the power of 10 is " -34" not 32

OpenStudy (anonymous):

Sorry..I had a random brain glitch.

OpenStudy (anonymous):

^better? xD

OpenStudy (anonymous):

Also it is 6.626 not 6.26

OpenStudy (anonymous):

Yikes. I'm really glitching. One sec. I'll fix it.

OpenStudy (anonymous):

E=hf=hc/λ E-energy of a photon h-plank's constant: 6.626*10^-34 m^2kg/sec λ-wavelength c-speed of light: 2.998*10^8 ms^-1 So basically use the E=hc/λ and then plug in E for E=hf Rearrange for E/h=f I think that's what doug is trying to get at.

OpenStudy (anonymous):

Perfect!! :)

OpenStudy (anonymous):

I'm like.....mixing Avogadro's number and Plank's constant....Dx gah...someone save me.

OpenStudy (anonymous):

Actually we can find the Wavelength associated with highest frequency in a simple way!!! E=hc/λ & E=hf There fore f = c/λ From this we can say that frequency is inversely proportional to wave length Hence for the lowest Wavelength we get the highest frequency

OpenStudy (anonymous):

*applauds*

OpenStudy (anonymous):

Thanks :) Did you get it @Grazes